OpenStudy (dallaslifebaby01):

What are the next three terms in the sequences below? 1.) 2, 5, 7, 12, ___, ___, ___ 2.) 9, 11, 20, 31, ___, ___, ___ 3.) 2, 5, 10, 50, ___, ___, ___ 4.) 1, 3, 3, 9, ___, ___, ___

OpenStudy (solomonzelman):

oh the patterns are: 2, 5, 7, 12, ___, ___, ___ 9, 11, 20, 31, ___, ___, ___ 2, 5, 10, 50, ___, ___, ___ 1, 3, 3, 9, ___, ___, ___ correct? (got a bit confused by the numbers in front)

OpenStudy (dallaslifebaby01):

yes let me edit it

OpenStudy (solomonzelman):

the last pattern is \(\Large\color{black}{ a_{n+1}=a_{n}\times a_{n-1} }\)

OpenStudy (solomonzelman):

the 3rd pattern is r=5/2 (geometric sequence)

OpenStudy (dallaslifebaby01):

i don't have anything to work it out on but im sitting some kids and I'm trying to go over their work

OpenStudy (solomonzelman):

the 2nd pattern is Fibonacci sequence, but they start from 9 and 11

OpenStudy (solomonzelman):

and the 1st sequence is also Fibonacci sequence, that starts from 2 and 5

OpenStudy (solomonzelman):

(i gave you the patterns)

OpenStudy (dallaslifebaby01):

okay thank you!